### v1

# DetailGen3D: Generative 3D Geometry Enhancement via Data-Dependent Flow Ken Deng1,2,3  Yuanchen Guo2  Jingxiang Sun1  Zixin Zou2  Yangguang Li2  Xin Cai4 Yanpei Cao2  Yebin Liu1  Ding Liang1,2

Abstract
Modern 3D generation methods can rapidly create shapes from sparse or single views, but their outputs often lack geometric detail due to computational constraints. We present DetailGen3D, a generative approach specifically designed to enhance these generated 3D shapes. Our key insight is to model the coarse-to-fine transformation directly through data-dependent flows in latent space, avoiding the computational overhead of large-scale 3D generative models. We introduce a token matching strategy that ensures accurate spatial correspondence during refinement, enabling local detail synthesis while preserving global structure. By carefully designing our training data to match the characteristics of synthesized coarse shapes, our method can effectively enhance shapes produced by various 3D generation and reconstruction approaches, from single-view to sparse multi-view inputs. Extensive experiments demonstrate that DetailGen3D achieves high-fidelity geometric detail synthesis while maintaining efficiency in training.
Figure 1: Our method demonstrates effective geometry refinement across various tasks and representations. In the images, the coarse geometry is displayed in gray  , while the refined geometry produced by our approach is shown in red  . On the right, zoomed-in details are provided to better observe the refinement effects.
1 Introduction
Obtaining high-quality 3D geometry has been a long-standing research focus in the fields of computer vision and graphics. High-quality 3D models are not only valuable in the film industry, video games, and virtual reality, but also play a crucial role in the rapidly advancing field of embodied intelligence, contributing significantly to simulation environments. Early approaches for high-quality multi-view stereo reconstruction rely on dense multi-view inputs [16, 14, 1, 94, 26, 7]. Although recent methods [75, 78, 25, 6] based on Neural Radiance Field (NeRF) [48] and 3D Gaussian Splatting (3DGS) [30] have improved the end-to-end performance, reconstructing high-quality geometry remains a challenging problem.
Recent methods for 3D generation from sparse or single views have evolved into two main paradigms: 1) optimization using 2D generative models and 2) direct training with 3D data. The first approach leverages pre-trained 2D vision models - notably, DreamFusion  [54] introduced Score Distillation Sampling (SDS) to align rendered views with text-conditioned distributions. Related methods  [79, 65] have extended this paradigm, though they consistently face geometric inconsistency issues, such as the “Janus problem” where details conflict across viewpoints. The second paradigm, training with 3D data, has seen multiple technical advances. Multi-view diffusion approaches  [40, 37, 36, 61, 33, 55, 44, 46, 60, 74] achieve view consistency but struggle with geometric coherence in fine details. Feed-forward Large Reconstruction Models  [24, 90, 32, 76, 106, 39, 81, 102] offer impressive speed but face resolution constraints. While recent 3D diffusion models  [99, 104, 103, 84, 56, 100] demonstrate quality improvements through direct 3D training, computational demands continue to limit their achievable resolution and detail fidelity.
While recent advances in 3D generation have shown promising results, the generated shapes often lack fine geometric details. Simply scaling up these generative models demands prohibitive computational resources. Traditional geometry refinement approaches that rely on high-resolution dense multi-view images are ill-suited for enhancing such generated shapes, as these additional inputs are typically unavailable. Furthermore, existing optimization-based methods that utilize normal or shading information [83, 97, 82, 87, 50] require precise texture alignment and struggle to preserve global shape coherence. To address these challenges, we propose DetailGen3D, a generative approach to 3D geometry refinement that learns to synthesize plausible fine-scale details directly from a collection of high-quality 3D shapes. By capturing the underlying geometric patterns through data-driven priors, our method can enhance coarse shapes while maintaining structural consistency, even in the presence of noise or incomplete information.
To tackle geometry refinement tasks with a generative model, a straightforward approach is to train a coarse geometry conditioned diffusion model, as seen in 2D image restoration methods [3, 96]. However, this method can be resource-intensive and may not be optimal for 3D tasks. Instead, inspired by Fischer et al. [12], who achieved effective image super-resolution without relying on large generative models, we propose a more training-efficient strategy. Rather than learning a complex mapping from noise to fine details, we focus on directly modeling the transformation between coarse and fine geometry using data-dependent rectified flow [41]. This approach utilizes optimal transport, which provides a direct and structured mapping between coarse and fine data points. By exploiting this coupling, we eliminate the need for the random coupling of noise and fine details, thus significantly reducing the training cost. This method is better suited for scaling up, offering a more efficient path to geometry refinement at larger scales.
Specifically, we introduce a training method called token matching. For the framework of our method, we follow 3DShape2Vecset [99]. In the 3D VAE encoding process, the query points obtained by downsampling are cross-attentioned with the global point cloud to get tokens. After denoising by DiT, the predicted tokens are sent to the 3D VAE decoder to obtain the implicit field as a 3D representation. In our method, after obtaining ground-truth (GT) geometry’s query points, we apply a nearest-neighbor algorithm to map the GT geometry’ query points across the entire coarse geometry point cloud. This approach ensures robust one-to-one token correspondence in space, maintains computational efficiency, and supports diverse geometric complexities.
Effective geometry refinement requires carefully designed training data. While high-quality 3D shapes are available, obtaining matching coarse-fine pairs poses a significant challenge. Simply applying traditional mesh simplification or smoothing operations often produces artifacts that differ fundamentally from the characteristics of generated shapes. To address this, we leverage an LRM-based model that reconstructs 3D geometry from sparse-view renderings of high-quality fine models. This approach naturally introduces neural network-induced variations in geometric details, closely mimicking the characteristics of shapes produced by generative models. These generated shapes naturally serve as coarse inputs for training.
In summary, our contributions are as follows:
  1. 1.
    We develop a novel generative geometry refinement algorithm, demonstrating its highly effectiveness for different geometric representations.
  2. 2.
    We propose a data-dependent rectified flow to incorporate coarse geometry information, enabling local distribution transformation from coarse to refined geometry.
  3. 3.
    We introduce a token matching training method that significantly enhances training efficiency and spatial correspondence accuracy.
2 Related Works
Optimization based Model Reconstruction. Optimization based multi-view reconstruction methods experienced a rapid development with the emergence of NeRF [48]. However, methods like NeRF and NeuS [75] rely on lengthy optimization processes and dense multi-view inputs to achieve detailed geometry. Although subsequent approaches aimed at reducing the number of required views [77, 4, 28, 31, 42, 45, 57, 73, 68, 92], exploring reflective object reconstruction [43, 71, 18], improving reconstruction speed [49, 63, 64, 5, 17, 78], have been introduced, their geometric performance remains relying on dense multi view as input. Following the development of 3DGS [30], a representation that present faster rendering speed but worse geometry quality, a series of works focused on geometric optimization have emerged, including  [20, 25, 47, 98, 6], yet they have not resolved the issues of geometry quality and reconstruction speed.
3D generation using 2D generative model. With the significant advancements in text-to-image generation models [58, 59, 2], methods for text to 3D generation based on SDS loss optimization  [54, 35, 88, 8, 79, 65, 69, 95, 34] have emerged, allowing the acquisition of 3D models without relying on multi-view inputs. However, these image diffusion models lack 3D priors, leading to the generation of 3D models that often suffer from the “Janus problem”, with poor alignment between geometry and texture, as well as substantial time overhead.
3D generation using 3D data training. Finetuning image diffusion models using Objaverse [9], Objaverse-XL [10] to generate reasonably consistent multi-views with 3D consistency, which could then be input into multi-view reconstruction models to obtain 3D geometry  [40, 37, 36, 61, 33, 55, 44, 46, 60, 74]. While methods combining these models with SDS optimization for 3D modeling are still relatively slow, their feedforward multi-view generation has significantly improved both generation quality and speed. However, optimization-based methods for obtaining 3D models still require several minutes to complete. Additionally, the generated results are still unsatisfactory because they are struggling to ensure 3D consistency in details. Some methods trying to improve multi-view consistency [70, 91, 80, 89], but geometry quality is even worse than previous methods since only 2D supervision is used. Using triplane as 3D representation, large-scale reconstruction model  [24, 32, 76, 90, 72, 66, 81, 106, 93] utilizing transformer architectures and triplane representations with fixed-view multi-view inputs, enabling reconstruction in just a few seconds. However, the extensive training time and GPU memory requirements of large reconstruction models severely limit their resolution. Generative models using triplane as 3D representation [19, 101, 23, 62, 85] meet the same problem as well. Using point cloud as 3D representation improves the efficiency [29, 51, 99, 104, 103] also brings high training costs. Some recent methods use voxel [56, 38] or gaussian [70, 102, 100] as 3D representation are also efficient.
3 Methods
Figure 2: (1) Inference pipeline. We use FPS-VAE to extract tokens of the coarse geometry generated or reconstructed, then input the coarse token and DINO feature [52] of the image prompt to DiT [53]. After denoising, we decode the predicted token using FPS-VAE decoder to obtain refined geometry. The inference process takes only a few seconds. (2) For training data, we use reconstruction results reconstructed by LRM using multi-views rendered from fine geometry as coarse geometry. (3) We demonstrate the token matching process on the left. On the right, for the top one, we only use part query points, which are located in quadrant one, and for the bottom one, we use full query points, which demonstrate tokens represent the space around the corresponding query points.
We formulate our geometry enhancement task as follows: given an input of coarse geometry obtained by reconstruction or generation and a corresponding view of the desired fine geometry, the objective is to produce fine geometry that aligns with both the image prompt and the input coarse geometry. The inference pipeline is shown in Fig 2.
3.1 Model Architecture
Our network structure mainly consists of two parts: FPS-VAE and Image conditioned DiT. FPS-VAE encoder is used to encode 3D shapes into tokens to save GPU memory. Different from  [103], which directly downsample full point cloud to obtain query points, we add farthest point sampling (FPS) [15] process after random downsampling of full point cloud and the decoder is used to decode tokens into the implicit field. The Image conditioned DiT is used to model the refinement process, using coarse geometry’s token as input and predicting fine geometry’s token.
FPS-VAE. The design of our FPS-VAE is primarily inspired by 3DShape2VecSet [99] and CLAY [103]. To improve detail capture, we increase the point sampling count from CLAY’s maximum of 8192 to 20480. Additionally, we introduce a two-step downsampling process for the point cloud: the first step uses random downsampling, followed by FPS [15] downsampling. This approach helps ensure that different tokens correspond to distinct spatial regions, thereby improving the model’s ability to capture fine-grained details.
Image conditioned DiT. Our DiT network is designed with efficiency in mind, comprising 24 transformer layers with a width of 768, 12 attention heads, and a latent length of 2048, totaling 368M parameters. To accommodate the relatively small width, we inject the timestep using the adaLN [27] method:
$z=z+g_{\text{msa}}\cdot\text{self\_attn}(\text{mod}(\text{LN}_{\text{msa}}(z),s_{\text{msa}},\gamma_{\text{msa}}))$ (1)
Turning to conditioning on the image prompt, we use cross-attention, ensuring that the image guides the refinement process effectively while maintaining robust feature capture. The conditioning is defined as:
$z=z+g_{\text{mca}}\cdot\text{cross\_attn}(\text{mod}(\text{LN}_{\text{mca}}(z),s_{\text{mca}},\gamma_{\text{mca}}),y)$ (2)
At last, tokens will pass through a feedforward network:
$z=z+g_{\text{ffn}}\cdot\text{FFN}(\text{mod}(\text{LN}_{\text{ffn}}(z),s_{\text{ffn}},\gamma_{\text{ffn}}))$ (3)
For encoding, we sample $N=20480$ surface points from a 3D mesh, which are then transformed into latent tokens $z$ using $\mathcal{E}(\cdot)$. Simultaneously, a pretrained ViT (DINO V2 [52]) processes the image prompt into a feature map $y$. The DiT then predicts the tokens for fine geometry, conditioned on these inputs, ensuring fine-grained detail capture during the refinement stage.
3.2 Data Dependent Rectified Flow
We use data-dependent rectified flow to model the distribution mapping of coarse and fine geometry due to its efficiency. The noise added to fine token is sampling from coarse distribution in the training stage and in the inference stage, the denoising process is begin with the sampling from coarse distribution. This modeling can natually almost predict coarse geometry without training. In contrast to cross attention which modeling the distribution mapping between noise and fine geometry, using coarse geometry as condition needs long time training to achieve noise to coarse mapping.
Rectified Flow and Loss Functions. Unlike previous generative models that model the mapping between a Gaussian distribution and the GT data distribution, we model the mapping between the coarse geometry distribution and the fine geometry distribution. Let $x_{1}$ represent the fine geometry and $x_{0}$ represent the coarse geometry. The geometry at time $t$, $x_{t}$, is given by:
$x_{t}=(1-t)\cdot x_{0}+t\cdot x_{1}$ (4)
The conditional probability $P(x_{1}|x_{t})$ is defined as:
$P(x_{1}|x_{t})=\mathcal{N}(x|(1-t)\cdot x_{0}+t\cdot x_{1},\sigma_{\min}^{2}\cdot I)$ (5)
The loss function $L(\theta)$ is:
$L(\theta)=\mathbb{E}_{t,x_{0},x_{1},y}\left[\left\|v_{\theta}(t,x_{t},y)-(x_{1}-x_{0})\right\|^{2}\right]$ (6)
where $v_{\theta}$ is parameterized by the DiT network discussed above with learnable parameters $\theta$, and $y$ is the image condition guiding the geometry refinement.
Noise Augmentation. The degraded geometric distribution is relatively independent so direct modeling of distribution transformation between coarse and fine geometry cannot well learn comprehensive refinement rules from the dataset, leading to noisy results. Inspired by [13], we apply noise augmentation to the token of coarse geometry, which helps improve the quality. During training, we apply noise to $x_{0}$ according to the DDPM [22] linear schedule at 400 timesteps. The noise $\epsilon$ is added to $x_{0}$ before obtaining $x_{t}$, following the DDPM [22] linear schedule. This is done because applying noise after obtaining $x_{t}$ would affect the efficiency of optimal transport, leading to an optimization path that is not straight. During inference stage, we also noise $x_{0}$ at 100 timesteps.
3.3 Token Matching
Due to the token representing the space area around the corresponding query points, we introduce a training method called token matching to match the input tokens with GT tokens as demonstrated in Fig 2. Without correct matching, the network not only learns the refine transformation but also location swapping because the tokens encoded by FPS-VAE are unordered.
Although attempts were made to match tokens in latent space, the correspondence is difficult to model effectively. In contrast, query points have explicit meaning. We explored the use of point cloud sorting algorithms, which work well for simple geometries but struggle with complex geometries due to their inability to preserve spatial relationships. Optimal Transport (OT) between coarse and fine query points can handle more complex shapes, but is computationally expensive and must be performed offline.
To balance effectiveness and efficiency, we propose a novel token matching method. First, we obtain the query points of fine geometry by downsampling its point cloud. Next, we apply a nearest-neighbour algorithm to identify the closest fine geometry query points in the coarse geometry point cloud, using these as the coarse geometry’s query points. This method is computationally efficient and operates online, maintaining model generalization and offering robust performance across both simple and complex geometries, while ensuring local refinement.
Finally, it is important to note that replacing query points with learnable queries as proposed in [99] in the VAE does not address the issue. Firstly, the VAE’s performance with learnable queries is inferior to that when using query points. Secondly, learnable queries prevent the guarantee that tokens with the same indices in the encoded tensors of the coarse and fine geometries will correspond to similar features, leading to inconsistencies in token interpretation.
3.4 Dataset Design
To achieve local transformation between coarse and fine geometries, well-aligned coarse-fine pairs are essential. For coarse-fine pair design, as shown in Fig 2, we directly utilized large-scale reconstruction model reconstruction result of multi-view of fine model as coarse model. This approach minimizes the gap between prior observations and practical application. Unlike optimization-based methods, which are computationally expensive and time-consuming, this model completes reconstructions within 20 seconds, significantly improving dataset creation efficiency. The coarse-fine pairs produced by this method are well-aligned in both shape and spatial coordinates.
Although some geometric degradation, such as surface noise, remains for simple objects, our method has the advantage of inherently replicating the artifacts typical of neural network reconstructions. These artifacts are difficult to reproduce using human-designed priors. As a result, this method provides high-quality coarse-fine pairs that are both practically aligned and consistent with the models that will be encountered during inference, ensuring effective learning of the local transformation process.
4 Experiment
Figure 3: We apply our method on input meshes reconstructed by different approaches (Instant3D [32], CRM [80], InstantMesh [89]).   represent coarse,   represent fine refinement results from our method. The top three objects are from Objaverse, while the bottom object is from GSO. More results can be found in the supplementary.
Figure 4: We apply our method on input meshes generated by different approaches (TripoSR [72], CRM [80], InstantMesh [89]).   represent coarse,   represent fine refinement results from our method. The top three objects are from Objaverse, while the bottom object is from GSO. More results can be found in the supplementary.
Our comprehensive evaluation of the model’s 3D geometric refinement capabilities is divided into three parts: 1). FeedForward-based Reconstruction: We tested the refinement ability of our method on the reconstruction results of various LRM models to demonstrate its effectiveness. 2). Generation: We assessed our method’s performance on the generation outputs from different LRM models, showcasing its versatility across diverse 3D representations. 3). Optimization-based Reconstruction: We applied our method to NeuS-like [75] optimization-based reconstruction, achieving impressive results that highlight its strong generalization. We also mixed results from different datasets, methods, and tasks for a user study, and conducted ablation studies on our design choices. See the supplementary for more details.
For the Objaverse [9] test set used, the IDs for models trained with different LRM methods are varied and not publicly available, which may lead to unfair comparisons between methods. However, we emphasize that our refinement model has never seen these 3D models during training, so its ability to refine the same object across different models fairly reflects the model’s generalization capability.
4.1 Feed-forward Reconstruction
Our task is to refine a given coarse geometry guided by images. To ensure that the coarse and refined geometries are spatially aligned to enable localized refinement, we use four GT views as input during training, obtaining a coarse model reconstructed by Instant3D [32]. Multi-view diffusion, while generating multiple views from a single image, can still introduce significant 3D inconsistencies, which severely affect reconstruction quality. In this experiment, we render the dataset according to the camera requirements of each LRM’s multi-view input, thus eliminating the impact of multi-view diffusion on geometry quality.
For the reconstruction task, we randomly sampled 199 models from the GSO [11] dataset and selected 350 unseen models from [9] as the evaluation dataset. We used Instant3D [32], CRM [80], InstantMesh [89] (testing both NeRF and Mesh, the checkpoints used in our work are all the official large versions provided.) as the reconstruction models for the coarse geometry. First, we rendered multi-view images of the GT according to the input viewpoint requirements of each large reconstruction model. This way, the reconstructed geometry can be assessed without the influence of different multi-view diffusion performances. We selected FID [21] as our evaluation metrics.
Visualization results can be found in Fig 3 and metrics results can be found in Tab 1. Our method fundamentally utilizes a generative model to add geometric detail to coarse geometry. Since only one view is used as guidance, this does not constrain the optimized geometry from being spatially consistent with the GT. Moreover, adding geometric detail without absolute spatial consistency with the GT can still increase loss. Therefore, we choose FID [21] as an evaluation metric. For the test images, we follow the setting of SDF-StyleGAN [105].
FID $\downarrow$ Reconstruction Generation
Coarse Fine Coarse Fine
Instant3D [32] 20.33 19.07 x x
TripoSR [72] x x 51.80 33.33
CRM [80] 40.13 25.29 48.91 33.74
InstantMesh(NeRF) [89] 58.35 29.51 50.12 25.75
InstantMesh(Mesh) [89] 35.32 24.45 33.85 25.79
Table 1: FID comparison of applying our method to different LRM results on reconstruction/generation task.
Figure 5: We apply our method on input meshes generated by different approaches (TripoSR [72], CRM [80], InstantMesh [89]) using GPTEval3D as input.   represent coarse,   represent fine refinement results from our method. More results can be found in the supplementary.
4.2 Generation
For the generation task, we randomly sampled 199 models from the GSO [11] dataset and selected 350 unseen models from Objaverse [9] as the evaluation dataset, same as mentioned before. Then used 110 images provided by GPTEval3D [86] as input, employing TripoSR [72], CRM [80] and InstantMesh [89] (testing both NeRF and Mesh, the checkpoints used in our work are all the official large versions provided.) as generation models for the coarse geometry. We selected FID [21] as our evaluation metrics. For the test images, we follow the setting of SDF-StyleGAN [105].
Although our method uses reconstruction results as input with GT as supervision, it can still generalize to non-reconstructed object, such as generated results. Our method shows outstanding performance in FID scores. As the visualization result illustrated in Fig 4 and metrics result shown in Tab 1, our method performs particularly well on complex objects, while previous methods lack robustness; with the same image input, they cannot guarantee consistently reasonable results across all generated models. Therefore, comparisons before and after refinement do not accurately reflect the capabilities of our model.
Compared to the reconstruction experiments, we conducted a more challenging test on the generation task. We used 110 images provided by GPTEval3D [86] as input. Due to the high level of difficulty, most single-view 3D generation models struggled to produce reasonable 3D models based on these images. We selected the plausible 3D models from this set and applied our refinement method. The results as illustrated in Fig 5 demonstrate the robustness of our approach to such challenging data.
4.3 Optimization-based Reconstruction
Figure 6: We apply our method on input meshes reconstructed by neus.   represent coarse,   represent fine refinement results from our method. The two objects on the left are from Objaverse, while the first object on the right is from GSO. More results can be found in supplementary.
Method Coarse w/o Token matching w/o Image condition w/o Noise augmentation Cross attention Sorting w/o training Our full method
FID $\downarrow$ 20.33 19.79 21.31 17.98 23.92 20.58 24.22 19.32
Table 2: FID scores evaluated on the Objaverse test set (350 objects) for different model ablations.
Figure 7: Visualization of ablation analysis on model designs.
In this experiment, due to the significant time cost of such methods, we were unable to test on the Objaverse [9] validation set (350 objects) and GSO [11] dataset (199 objects) as in the previous two experiments. Instead, we selected complex objects for reconstruction. For Neus [75], to further improve speed, we used Instant-NSR [17] implementation. For the multi-view data, we rendered a uniformly distributed set of 40 views as input.
Optimization-based reconstruction algorithms produce satisfactory details in multi-view reconstructions. However, due to the lack of prior guidance, they are prone to artifacts in scenarios with complex occlusions, and the smoothness of the surface remains imperfect. The refinement results as illustrated in Fig 6 demonstrate that our method successfully refines and corrects irregular geometric details.
4.4 User Study
Existing metrics do not fully capture the refinement level we aim to measure. We evaluated this by presenting four views (azimuths: 45°, 135°, 225°, 315°; elevation: 15°) of both coarse and refined geometries, rendered with consistent code. Images were shuffled, and volunteers voted on which geometry better matched the reference. Results show 94% of votes favored the geometry refined by our method.
4.5 Ablation Study
In our ablation study, we tested the effects of removing token matching, using cross attention to inject coarse geometry information, removing image prompts, removing noise augmentation, and using a point cloud sorting algorithm to query point cloud, inference without training for reference. We trained for approximately 100,000 steps on a dataset of 40,000 objects (subsample from the training set), with the results evaluated on a test set of 350 objects from Objaverse (same as in previous experiments). We used FID as the evaluation metric on the rendered images following the SDF-StyleGAN [105] settings. The FID results are shown in Tab 2, and the visualizations are presented in Fig 7.
Experimental results indicate that removing token matching slows convergence, as the model must simultaneously learn both refinement transformations and location swapping; it is hard to generate unseen details in the coarse geometry without the image condition; cross-attention modeling lacks the constraint of one-to-one token correspondence between coarse and fine models, making the noise-to-fine process inefficient and yielding worse geometric quality after refinement; applying point cloud sorting to the query point cloud yields no clear benefit, and inference without training is almost as effective in predicting coarse geometry.
Removing noise augmentation results in the lowest FID, as shown in Tab 2. We speculate that this is primarily due to two reasons: first, the Inception model [67], which computes FID, is not sensitive to the noise perception of the ensemble model, and second, the image conditioning ensures the results more consistent with the ground truth, even without noise augmentation. While these factors contribute to low FID, the visual degradation remains clearly noticeable, as demonstrated in Fig 7 and further illustrated in the supplementary material.
5 Discussion and Conclusion
Limitation and future work. As a generative model using a single image prompt, our approach cannot guarantee refinement in the exact direction of the ground truth. Future work will focus on enabling multi-view guided refinement to capture global information and improve alignment with the ground truth.
Conclusion. In this paper, we present a generative 3D geometry refinement method using an image as a prompt. We introduce a training technique called Token Matching for localized geometry refinement. This approach proves highly effective in both reconstruction and generation tasks, delivering refined results across diverse datasets, particularly for complex geometries.